OpenStudy (bananas1234):

Find the degree of the polynomial. 4p^2 – 2pq + 3q^2 1 2 3 4

OpenStudy (zzr0ck3r):

The degree of each term in a polynomial in two variables is the sum of the exponents in each term and the degree of the polynomial is the largest such sum.

OpenStudy (welshfella):

the degree is the highest power in the expression p^2 is degree 2 what is the degrees of the other terms?

OpenStudy (zzr0ck3r):

they all sum to the same number

OpenStudy (zzr0ck3r):

this is not the definition of degree to variables @welshfella

OpenStudy (zzr0ck3r):

two*

OpenStudy (bananas1234):

hmm ok im still a bit confused

OpenStudy (welshfella):

2pq is degree 2 right?

OpenStudy (welshfella):

p^1 q^1

OpenStudy (welshfella):

1+1 = 2

OpenStudy (bananas1234):

That makes sense @welshfella Is that correct to you @zzr0ck3r ?

OpenStudy (zzr0ck3r):

right...so this is a polynomial in two variables p, q (unless one is supposed to be constant) the degree of each term of the polynomial is the degree of the two variables in that term added together, so for the first term you have p^2q^0 and 2+0 =2 in the second term you have p^1q^1 and 1+1 = 2 and finally in the last term we have p^0q^2 and again 2+0 = 2, so we have degrees 2, 2, and 2. we pick the largest of these for the degree of the polynomial (here it is obviously 2).

OpenStudy (bananas1234):

Thank you i think i get it now! @zzr0ck3r